EMPIRICAL LOWER STATE ENERGIES OF ^{13}CH_{4} AT 1.66 μm USING 296 K AND 81 K SPECTRA

O. M. LYULIN, Laboratory of Theoretical Spectroscopy, Institute of Atmospheric Optics, Siberian Branch, Russian Academy of Sciences, 1, Akademicheskii av.,634055 Tomsk; S. KASSI, A. CAM-PARGUE, Laboratoire de Spectrométrie Physique (associated with CNRS, UMR 5588), Université Joseph Fourier de Grenoble, B.P. 87, 38402 Saint-Martin-d’Hères Cedex, France; K. SUNG, \underline{L. R. BROWN}, Jet Propulsion Laboratory, California Institute of Technology, 4800 Oak Grove Dr.,Pasadena, CA 91109, U.S.A..

The high resolution absorption spectra of ^{13}CH_{4} were recorded at 81 K by differential absorption spectroscopy using a cryogenic cell and a series of Distributed Feed Back (DFB) diode lasers at room temperature by Fourier transform spectroscopy (a Bruker IFS-125HR at JPL). Empirical line lists were constructed containing, respectively, 1629 ^{13}CH_{4} transitions detected at 81 K (5852 - 6124 cm^{−1}) and 3481 features measured at room temperature (5850 - 6150 cm^{−1}); the minimum observed intensities were, respectively, 3×10^{−26} and 4×10^{−25} cm/molecule at 81 K and 296 K. From the variation of the cold and room temperaure line intensities, empirical lower state energies were derived for 1196 ^{13}CH_{4} transitions. Over 400 additional weak features, detected at 81 K, could not be matched to lines observed at room temperature. The observed intensities represent 99.2% and 84.6% of the total absorbance at 81 K and 296 K, respectively. The quality of the resulting empirical low energy values is demonstrated by the excellent agreement with the already-assigned transitions and the clear propensity of the empirical low J values to be close to integers.^{a}

aPart of the research described in this paper was performed the Jet Propulsion Laboratory, California Institute of Technology, under contracts and cooperative agreements with the National Aeronautics and Space Administration.